How many grams of each product result from the following reac-

tions, and how many grams of which reactant are left over?

(a) 11.3 g NaCl2 + 13.5 g AgNO32 ¡1x g AgCl2 + 1y g NaNO32

(b) 12.65 g BaCl22 + 16.78 g H2SO42 ¡1x g BaSO42 + 1y g HCl2


1
Expert's answer
2021-01-05T05:57:55-0500

A. NaCl + AgNO3 ---> AgCl + NaNO3

Molar mass of NaCl= 23+35.5= 58.5g

Molar mass of AgNO3= 108 + 14 + 48 = 170g

Molar mass of AgCl= 108+35.5= 143.5g

Molar mass of NaNO3= 23+14+48= 85g

58.5g of NaCl react with 170g of AgNO3

11.3g of NaCl will react with 32.84g of AgNO3

Since only 13.5g of AgNO3 is present, it is the limiting reagent.

For AgCl:

170g of AgNO3 yields 143.5g of AgCl

13.5g of AgNO3will yield 143.5x13.5/170 = 11.4g of AgCl


For NaNO3:

170g of AgNO3 yields 85g of NaNO3

13.5g of AgNO3 will yield 85x13.5/170 = 6.75g of AgCl


Since 13.5g of AgNO3 reacted,

170g of AgNO3 reacts with 58.5g of NaCl

13.5g of AgNO3 will react with 58.5x13.5/170 = 4.65g of NaCl

Amount of NaCl left = 11.3-4.65= 6.65g of NaCl


B. BaCl2 + H2SO4 -----> BaSO4 + 2HCl

Molar masses:

BaCl2= 208.33g

H2SO4= 98g

BaSO4= 233.33g

HCl= 36.5g

98g of H2SO4 react with 208.33g of BaCl2

16.78g of H2SO4 will react with 35.67g

Since only 12.65g of BaCl2 ia present, it is the limiting reagent.


For BaSO4:

208.33g of BaCl2 yields 233.33g of BaSO4

12.65g of BaCl2 will yield 233.33x12.65/208.33 = 14.17g

For HCl:

208.33g of BaCl2 yields 2(36.5)g of HCl

12.65g of BaCl2 will yield 2(36.5)x12.65/208.33 = 4.43g of HCl


Since only 12.65g of BaCl2 is available.

208.33g of BaCl2 react with 98g of H2SO4

12.65g of BaCl2 will react with 98x12.65/208.33 = 5.95g

Amount of H2SO4 left = 16.78-5.95 = 10.83g